# SECURA: Sigmoid-Enhanced CUR Decomposition with Uninterrupted Retention and Low-Rank Adaptation in Large Language Models

### v1

@@ 3 Methodology
Our proposed SECURA method is based on three key components: CABR-LoRA reverse decomposition fine-tuning, the SigNorm Normalization algorithm, and two types of SECURA Merge method. Drawing inspiration from dynamic network pruning and CUR-decomposition, we hypothesize that parameters with higher weights in the original model are more important, as they store more critical knowledge. By constraining these important parameters and focusing on fine-tuning the less important ones, SECURA can achieve remarkable performance while significantly mitigating catastrophic forgetting. Appendix B shows all intuitive equation derivation.
@@
3.1 CABR-LoRA Reverse Decomposition Initialization
As shown in Figure 1, we initialized CABR-LoRA as a reverse LoRA decomposition of the $U$ matrix. This allows the "not important" columns and rows to remain smaller while maintaining a higher-dimensional space for fine-tuning SECURA. The formula for initialize of SECURA is:
@@ 3.1 CABR-LoRA Reverse Decomposition Initialization
CABR-LoRA is based on CUR-LoRA, utilizing CUR decomposition to break down the weight matrices. CUR-LoRA reduces the number of trainable parameters to locate lest important $\boldsymbol{C}$ and $\boldsymbol{R}$, but this often leads to the loss of important parameters, causing performance degradation. To address this, we introduce CABR decomposition, which decomposes the $U$ matrix, into two matrices: $\mathcal{W}_{\text{A}}$ and $\mathcal{W}_{\text{B}}$, which shapes are $r\times m$ and $m\times r$. Respectively here $m>r$. $\mathcal{W}_{\text{A}}$ is initialized via SVD decomposition, while $\mathcal{W}_{\text{B}}$ is initialized with zeros. The decomposition follows the formulas below:
@@
By initialize $\mathcal{W}_{A}$ with SVD decomposition, CABR-LoRA preserved the most knowledge learned before. The initialization of $\mathcal{W}_{B}$ with zeros ensures that this matrix will fine-tune from the base LLM without altering the critical knowledge.
@@ 3.2 SigNorm Normalization Algorithm
The CABR decomposition provides a higher dimension of fine-tuning, but only using CABR-LoRA alone does not fully address catastrophic forgetting, which is crucial for performance in lifelong learning. To tackle this, we introduce the SigNorm Normalization algorithm, inspired by dynamic network pruning. This algorithm adjusts the weight updates by controlling the change magnitude between the former and new weights, provide a stable parameter update figure 2.
The core idea is to apply a Sigmoid function to dynamically adjust the magnitude of weight updates by getting a weight pruning matrix which value between range in $[1,2]$. Because sigmoid function mainly focus on central value’s change and suppresses outliers on both ends. We can allow SECURA deal with whole matrix’s changing tendency while limiting the extreme parameters’ effect. At the same time, restricting the update amplitude of lower-norm columns and rows ensures that important parameters(with higher norms)’ magnitude experience minimal change, while less important parameters (with smaller norms) undergo more significant updates. The former weight matrix is denoted as $W_{\text{former}}$, and the SECURA’s matrices as $A$ and $B$.
First, the merged weight matrix is computed as:
@@
Next, we calculate the MagLossMatrix $\mathcal{M}_{\text{Mag}}$, which captures the relative magnitude change between the new and former weights. Here $\epsilon$ is a small value to avoid division by too small parameters:
@@
Then, we normalize the MagLossMatrix to the range $[-0.5,0.5]$ to balance the values on both sides of zero:
@@
Once we have weight pruning matrix $\mathcal{M}_{\text{Res}}$ the megred matrix $W_{\text{SECURAMerged}}$, will adjust the update process by dividing the SigNorm pruning-weight $\mathcal{M}_{\text{Res}}$:
@@
Figure  3 shows, this dynamic update process allows controlled pruning of weights, where updates for important parameters are slowed down and less important one remains, helping preserve important knowledge and mitigating catastrophic forgetting.
@@ 3.3 Merge with Basic Model:
M1 offers better fine-tuning but sacrifices some knowledge retention, while M2 retains more knowledge but sacrifices fine-tuning performance due to shallower merging. Based on our ablation study 4, for tasks focused on knowledge retention, we recommend M2, and for tasks prioritizing fine-tuning performance, M1 is preferred. M1 also provides a balance, maintaining a high level of knowledge retention while delivering better fine-tuning results.
@@ 4.1 Experimental Setup
  • CUR-LoRA (Fawi, 2024):CUR-Decomposed low-rank adaptation
@@ Appendix D Experimental hyper parameters on three main experiments
Table 123,  4 and Figure 4 show all hyper-parameters among different LoRA baselines and SECURA using in three main experiments.
@@
Exp 3
Rank 8 8 - 8 -
learning rates 1e-5 1e-5 1e-5 1e-5 1e-5
CABR/CUR r Rank - - 150 - 150
CABR m Rank - - - - 200
Fusion Step 2000 2000 2000 2000 1
@@ Appendix E Experimental Results on All Datasets
Table 789,  11 and  10 show all results among different LoRA baselines and SECURA using Gemma2 2b, Qwen2 1.5B, Qwen2 7B, LLaMA3 7B and LLaMA3.1 8B backbones.

### v3

@@ 3 Methodology
Our proposed SECURA method is based on three key components: CABR-LoRA inverse decomposition fine-tuning, the SigNorm Normalization algorithm, and two types of SECURA Merge method. Drawing inspiration from dynamic network pruning and CUR-decomposition, we hypothesize that parameters with higher norms in the original model are more important, as they store more critical knowledge. By constraining these important parameters and focusing on fine-tuning the less important ones, SECURA can achieve remarkable performance while significantly mitigating catastrophic forgetting. Appendix B shows all intuitive equation derivation.
@@
3.1 CABR-LoRA Inverse Decomposition Initialization
CABR-LoRA is based on CUR-LoRA, utilizing CUR decomposition to break down the weight matrices. As shown in Figure 1, CABR-LoRA is initialized using a inverse low-rank decomposition of the $U$ matrix. This allows the "less important" columns and rows to remain smaller while maintaining a higher-dimensional space for fine-tuning SECURA. The formula for initialize of SECURA is:
@@ 3.1 CABR-LoRA Inverse Decomposition Initialization
Where $\mathcal{W}_{\text{CABRWeight}}$ is the initialized CABR-LoRA weight, $C$ and $R$ represent the less important columns and rows, while $A_{\text{initial}}$ and $B_{\text{zeros}}$ are the trainable matrices with weights $\mathcal{W}_{\text{A}}$ and $\mathcal{W}_{\text{B}}$.
CUR-LoRA reduces the number of trainable parameters to locate lest important $\boldsymbol{C}$ and $\boldsymbol{R}$, but this often leads to the loss of important parameters, causing performance degradation. To address this, we introduce CABR decomposition, which decomposes the $U$ matrix, into two matrices: $\mathcal{W}_{\text{A}}$ and $\mathcal{W}_{\text{B}}$, which shapes are $r\times m$ and $m\times r$. Respectively here $m>r$. $\mathcal{W}_{\text{A}}$ is initialized via SVD decomposition, while $\mathcal{W}_{\text{B}}$ is initialized with zeros. By initializing $\mathcal{W}_{A}$ with SVD decomposition, CABR-LoRA effectively preserves prior knowledge. The initialization of $\mathcal{W}_{B}$ with zeros ensures that this matrix will fine-tune from the base LLM without altering the critical knowledge. The decomposition follows the formulas below:
@@ 3.2 SigNorm Normalization Algorithm
While CABR decomposition provides a higher dimension of fine-tuning, only using CABR-LoRA alone does not fully address catastrophic forgetting, which is crucial for performance in lifelong learning. To tackle this, we introduce the SigNorm Normalization algorithm, inspired by dynamic network pruning. This algorithm adjusts the weight updates by controlling the change magnitude between the former and new weights, provide a stable parameter update (figure 2).
The core idea is to employ a Sigmoid-based dynamic scaling mechanism to regulate weight updates. This method constructs a weight pruning matrix whose values are constrained within the range $[1,2]$. The Sigmoid function primarily emphasizes central value adjustments, effectively suppressing extreme variations. We can allow SECURA deal with whole matrix’s changing tendency while limiting the extreme parameters’ effect. At the same time, restricting the update amplitude of lower-norm columns and rows ensures that important parameters(with higher norms)’ magnitude experience minimal change, while less important parameters (with smaller norms) undergo more significant updates. The former weight matrix is denoted as $W_{\text{former}}$, and the SECURA’s matrices as $A$ and $B$.
The merged weight matrix is computed as:
@@
Next, we calculate the Magnitude Loss Matrix $\mathcal{M}_{\text{Mag}}$, which captures the relative magnitude change between the new and former weights. Here $\epsilon$ is a small value to avoid division by too small parameters:
@@
Then, we normalize the $\mathcal{M}_{\text{Mag}}$ to the range $[-0.5,0.5]$ to balance the values on both sides of zero:
@@
Once we have weight pruning matrix $\mathcal{M}_{\text{Res}}$ , the megred matrix $W_{\text{SECURAMerged}}$, will then adjust the update process by dividing the SigNorm pruning-weight matrix $\mathcal{M}_{\text{Res}}$:
@@
This dynamic update process allows controlled pruning of weights, where updates for important parameters are slowed down and less important one remains, helping preserve important knowledge and mitigating catastrophic forgetting(Figure  3).
@@ 3.3 Merge with Basic Model:
For the two different merging methods, M1 achieves better fine-tuning performance but at the cost of some knowledge retention, while M2 preserves more knowledge but sacrifices fine-tuning effectiveness due to its shallower merging. Based on our ablation study (4), we recommend M2 for tasks emphasizing knowledge retention (e.g., medical and economics tasks) and M1 for tasks prioritizing fine-tuning performance (e.g., mathematics and programming tasks). Overall, M1 strikes a better balance, maintaining strong knowledge retention while achieving superior fine-tuning results.
@@ 4.1 Experimental Setup
  • CUR-LoRA (Fawi, 2024): CUR-Decomposed low-rank adaptation.
@@ Appendix D Experimental hyper parameters on three main experiments
Table  6 and  7 show all hyper-parameters among different LoRA baselines and SECURA using in three main experiments.
@@
Table 6: hyper-parameters of all experiments
Exp 3 SEQ EWC ER ILoRA CUR-LoRA SECURA
Rank - - - 8 - -
Learning Rates 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5
CABR/CUR r Rank - - - - 150 -
CABR m Rank - - - - 200 -
Fusion Step 2000 2000 2000 2000 1 1
Trainable Parameters (each layer) FT FT FT 2.62e+5 1.20e+5 1.00e+5
@@ Appendix E Experimental Results on All Datasets
Table 8910,  12 and  11 show all results among different LoRA baselines and SECURA using Gemma2 2b, Qwen2 1.5B, Qwen2 7B, LLaMA3 7B and LLaMA3.1 8B backbones.